OpenStudy (anonymous):

this i dont know how to start... integrate ln(x^4)/x dx

OpenStudy (anonymous):

The derivative of ln x is 1/x; you should be able to use U substitute to figure out the rest.

OpenStudy (anonymous):

int by parts?

OpenStudy (anonymous):

start with \[4\int \frac{\ln(x)}{x}dx\] the u - sub

OpenStudy (anonymous):

ohh no dimple integration

OpenStudy (anonymous):

Start by rewriting ln(x^4) as 4ln(x) Then use ln(x)=u (1/x)dx=du Then you have: \[4\int\limits udu=2u^2+C=2\ln^2(x)+C\]

OpenStudy (anonymous):

what malevolence said

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

My answer of course is assuming that x>0

OpenStudy (anonymous):

:)